# Robustness of Decentralised Learning to Nodes and Data Disruption

### v1

@@ 4.4 Network and data resilience scenarios
A node in the network can contribute to the decentralised learning process in two ways. It can be well positioned in the network facilitating knowledge (i.e., model) circulation and/or can be endowed with training data to be used for model training and update. Thus, a disruption might have a different effect depending on whether the switched-off nodes only facilitate model dissemination or also possess local training data on which the local model can be trained. For this reason, we will consider different scenarios. First (Case 1), we investigate network resilience when the disrupted nodes are not assigned local data. Thus, the highly central nodes (those that are cut off) will act as bridges but will not contribute knowledge to the learning process. They will pass along aggregate models without using their local data for training because they don’t have any. Next, we consider a situation where the highly central nodes are assigned data, train locally on their own data set, and share their updates with their neighbours. Thus, they are no longer just passive elements of the network. In these settings, the characteristics of the data items assigned to the disrupted nodes may play a crucial role in the resulting learning performance. Thus, we consider two different data distributions. We refer to Case 2 as the configuration in which disrupted nodes are assigned data in an IID fashion with respect to surviving nodes. In practice, this means that we assign exactly the same number of data samples per class to all nodes, regardless of whether they survive or don’t survive the disruption. Finally, in Case 3 we focus on a non-IID data distribution as follows. The available class labels are split into two groups, $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$: nodes that will be disrupted get disproportionately more data in the second label group than the other nodes, while all nodes get the same number of data for the first label group. With this configuration, we are testing what happens when disrupted nodes contribute much more knowledge to the training process than the other nodes. This is because larger local datasets can lead to better-trained models, thus enabling more knowledge extraction.
@@
For what concerns the training dataset, for our experiments, we employ the widely used MNIST image dataset [20]. This dataset contains a set of handwritten digits; thus, data are divided into 10 classes (for digits from 0 to 9). The collaborative task undertaken by the nodes is then a standard image classification problem. This problem is relevant to a disaster scenario, where individuals within the affected community can leverage their smartphones or other devices with built-in cameras to capture images of the affected areas, including collapsed buildings, road blockages, or other hazardous conditions. The MNIST dataset contains 60,000 training images (approximately 6,000 per class) and 10,000 test images (1,000 per class). We evaluate the impact of the disconnection of central nodes by measuring the accuracy obtained by the local models on the standard MNIST test dataset over time. Note that all classes are equally represented in the test set and that the test set is common to all nodes.
@@
Classification on the MNIST dataset is a relatively easy learning task. Thus, in order to stress test our decentralised learning process and to capture a realistic scenario [^3] We anticipate that in scenarios where nodes represent user devices collaborating on decentralised learning tasks, the local datasets will be relatively small, as users actively collect data samples themselves., we assign very few images per class to each node. In Case 1, we distribute 7 images per class to non-disrupted nodes, leading to a local training dataset of 70 images and 6,300 images overall in the network. Recall, in fact, that in Case 1 the 10 disrupted nodes do not own any local data, so the remaining 90 nodes are the only ones assigned data. In Case 2, we keep the same IID data distribution, but this time, we involve disrupted nodes. In order to get approximately the same dataset size distributed across the whole network, this time, we assign 6 images per class to each node, leading to a local training dataset of 60 images and, globally, 6,000 images in the network (given that images cannot be fractional, this is the best trade-off to make Case 1 and Case 2 comparable). In Case 3, recall that class labels are split into two sets $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$. Images belonging to $\mathcal{L}_{1}$ are distributed in an IID fashion across all nodes by simply splitting the approximately 6,000 images among all nodes (hence, each node gets around 60 images per class). Instead, only a small number of images belonging to $\mathcal{L}_{2}$ are assigned to nodes in $\mathcal{V}_{s}$ (we test configurations with 10, 20, and 30 images per class), while all the others are split equally among disrupted nodes (those in $\mathcal{V}_{d}$), which end up with approximately 500 images per $\mathcal{L}_{2}$-class. In Case 3, disrupted nodes are disproportionately better at classifying $\mathcal{L}_{2}$-images, because they enjoy a rather larger training set with respect to surviving nodes. Thus, with this configuration, we are able to assess how well the surviving nodes can retain the knowledge of disrupted nodes once the latter ones are cut off from the network. The three different scenarios and the corresponding data distributions are summarised in Table 1.
@@
Case 3 Classes [0-4], $\sim 60$ images per class Classes [0-4], $\sim 60$ images per class
Classes [5-9], $\sim 500$ images per class Classes [5-9], 10/20/30 images per class
Table 1: Summary table for the data distributions in the three considered scenarios.
4.5 Other settings
@@ 4.6 Evaluation metrics
Here, we introduce the metrics used to evaluate the robustness of decentralised learning to node disruption. A direct measure of classification performance is the accuracy achieved by nodes using their local models [^4] Note that local models here are the result of progressively averaging models obtained from neighbours. Thus they embed global knowledge. and applying it to the test set, which is common to all nodes. Given that disrupted nodes do not participate in the learning any more after disruption, we do not consider the accuracy they achieve in our analyses. To obtain a compact view of the performance, we also consider the average accuracy across non-disrupted nodes, which is computed as follows.
@@
The mean accuracy of the system is defined as the mean over all the $\mathcal{V}_{s}$ nodes:
$\overline{A}(t)=\sum_{i\,\in\,\mathcal{V}_{s}}\frac{1}{|\mathcal{V}_{s}|}A_{i}(t).$ (2)
$\overline{A}(t)$ will be used to measure the overall system’s performance.
When the top 10% central nodes are disrupted, the connectivity in the graph drastically decreases. As illustrated in Figure 2(b), many nodes become completely disconnected, while others form smaller clusters. Reduced connectivity results in fewer opportunities for communication and collaboration, impairing the decentralised learning process. It is anticipated that nodes losing more connections, and particularly those becoming isolated, will be affected most severely. Thus, we also define the mean accuracy for the individual clusters as follows.
Definition 4.2 (Mean accuracy within clusters).
The mean accuracy over clusters of size $c$ is defined as the mean over all $\mathcal{V}_{s}$ nodes belonging to the clusters of the chosen size $c$.
@@
where $\mathcal{C}_{c}$ denotes the set of nodes belonging to a cluster of size $c$.
The mean accuracy over clusters will be used to measure a more fine-grained impact of the disruption on the system, accounting for localised effects. Note that we aggregate over all nodes belonging to clusters with size $c$, regardless of whether they are in the same cluster. In the case of clusters of size 1, for example, this allows us to consider all isolated nodes together.
@@ 5 Results
In this section, we discuss the impact of central nodes being cut off from the network due to a disruption, and we focus on the three scenarios, Case 1, Case 2 and Case 3, illustrated in the previous section. Recall that in Case 1 and Case 2, nodes have an IID distribution of local data, with the difference that in Case 1, highly central nodes do not have local data, and thus they contribute only by connecting elements in the graph topology. The data that these nodes hold locally in Case 2 are redistributed in Case 1 among the other nodes, preserving the IID property among them to maintain the total number of data points equal at the overall network level. Case 3 is different from the other two, as we use it to test the impact of disruption when the data distribution across nodes is non-IID, and specifically when the disrupted nodes have much more local data (hence knowledge) than the others. First, in Sections 5.1-5.4, we discuss and compare Case 1 and Case 2, as in both the disrupted nodes do not enjoy more knowledge than the other nodes. Then, Case 3, in which disrupted nodes have much more knowledge than the other nodes, will be discussed in Section 5.5. Due to the significantly different data distributions between Cases 1-2 and Case 3, a direct quantitative comparison would not be appropriate.
@@ 5.1 Baseline performance when no disruption occurs
Figure 3 showcases what we already anticipated in Section 4.3: when two learning processes have differing speeds, a disruption occurring at a fixed time $t$ will encounter two systems at distinct stages of the learning process. While this scenario realistically captures the notion that faster systems are more likely to have acquired sufficient knowledge when disrupted, comparisons between the two systems for what-if analyses may be less informative. The discrepancy between the two scenarios is depicted in Table 2, which illustrates the communication round at which the mean accuracy of the system reaches the specified accuracy threshold in both cases. As expected, Case 2 reaches the accuracy threshold faster than Case 1. Therefore, as discussed in Section 4.3, in this study, we chose to assess the robustness of the two systems when disruptions occur at the same average accuracy level for both. The objective is to compare the consequences of disruption in systems that had achieved similar skill levels in terms of learning.
@@ 5.2 Impact of the time of disruption
Figure 4: Accuracy curves for each surviving node in Case 1 (a) and Case 2 (b). From left to right: increasing accuracy threshold condition for the disruption. The curves are coloured based on the size of the cluster to which the corresponding node belongs after the disruption. Accuracy threshold 75% is omitted for ease of visualisation, as it only showed an intermediate behaviour between threshold 70% and 80%.
@@
In Figure 4, we show the accuracy of all the surviving (i.e., non-switched-off) nodes in Case 1 and Case 2 when the disruption happens at different accuracy thresholds. The colouring here is defined based on the size of the cluster the nodes belong to after the disruption occurs. The clusters are the different connected components left after the central nodes are switched off due to the disruption, as discussed in Section 4.2.
@@
First, we focus on the effect of different disruption times in Figure 4. The later the disruption happens, i.e., going from left to right, the higher the mean accuracy of the system, as more and more nodes have been able to reach a higher level of accuracy before the cut-off, resulting in a narrower curve beam. This result holds for both scenarios and follows naturally after noticing that the central nodes are able to connect the network and, in Case 2, share their information for more communication rounds before getting switched off as the accuracy threshold increases. In all cases, after the disruption occurs, most curves stop improving and tend to cluster into different groups of similarly performing nodes. As the curve colours show, these groups reflect the cluster size to which the nodes belong. Recall that since the clusters’ sizes are calculated after the removal of the special nodes, nodes having cluster size 1 are isolated nodes. As we can see in Figure 4, isolated nodes have the lowest values of accuracies, and the accuracy generally increases with the increasing value of the cluster size. This comes naturally from the fact that more connected nodes enjoy more information flow due to collaborative learning, thus increasing their performance. Furthermore, by examining the curves for the two largest clusters (size 25 and size 29), we can observe that the larger the cluster size, the more a node’s performance becomes independent of the disruption time. This indicates that the cluster has sufficient data to compensate for the disruption’s effects, no matter when it occurs. Conversely, we notice a strong correlation between the accuracy level and the timing of disruptions for the isolated nodes: the later the disruption occurs, the better the accuracy level that the isolated nodes can achieve. Naturally, this is because they have longer access to more information. Clusters of intermediate sizes fall within these two boundary conditions.
@@
Interestingly, in Figure 4, we note that for later disruption times, isolated nodes, as well as nodes belonging to small clusters, reduce the accuracy they achieve with respect to the one they had achieved at the disruption time. This effect is not present for nodes in larger clusters. The intuition is that large clusters can compensate for the lack of connectivity thanks to the overall set of data residing at the nodes in the cluster. If the cluster is large enough, the effect of disruption can be compensated reasonably well. Otherwise, nodes in smaller clusters are not able to recover knowledge that is lost as a side effect of disruption. We analyse this effect in more detail in the next section.
@@ 5.3 Connectivity loss vs connectivity and data loss
Figure 6: Mean local accuracy difference for the largest cluster (left) and the isolated nodes (right).
@@
This is further confirmed at a more granular level when we compute the same difference but focusing on the largest cluster (Figure 6(a)) and the isolated nodes (Figure 6(b)), thus computing the difference between the within-clusters accuracy, as per Equation 3. The behaviour of the large cluster is similar to that observed for the system overall. Specifically, we note that (i) the condition of Case 1 is more favourable than that of Case 2 and that (2) after sufficient time, the accuracy threshold at which disruption occurred does not differentiate the two conditions. Moreover, the case of isolated nodes shows that when disruption occurs at a (relatively) low accuracy, the presence of even a small amount of additional data locally allows the model of Case 1 to reach a higher accuracy than the model trained in Case 2. However, this difference vanishes when disruption occurs at a high accuracy threshold.
All in all, our results suggest that having more data available “somewhere” in the network (as in Case 1) is more beneficial than concentrating the same amount of data on central nodes, if those nodes are at risk of disruption (as in Case 2), while the main advantage of the latter configuration is limited to a speed-up of the decentralised learning process if no disruption occurs (as shown by Figure 3). This confirms that, as long as data are present in the network, decentralised learning is able to exploit them through collaborative training, even in the presence of significant disruptions to the network structure.

### v2

@@ 4.3 Network and data resilience scenarios
A node in the network can contribute to the decentralised learning process in two ways. It can be well positioned in the network facilitating knowledge (i.e., model) circulation and/or can be endowed with training data to be used for model training and update. Thus, a disruption might have a different effect depending on whether the switched-off nodes only facilitate model dissemination or also possess local training data on which the local model can be trained. For this reason, we will consider different scenarios: one in which disrupted nodes are not assigned data and two in which they are. In the latter, the highly central nodes have local data, train locally on their own data set, and share their updates with their neighbours. Thus, they are no longer just passive elements of the network. In these settings, the characteristics of the data items assigned to the disrupted nodes may play a crucial role in the resulting learning performance. For this reason, we test different data assignments to disrupted nodes.
Case 1: Disrupted nodes are not assigned local data
When the disrupted nodes are not assigned local data, the highly central nodes (those cut off) will act as bridges but not contribute knowledge to the learning process. They will pass along aggregate models without using their local data for training because they don’t have any.
Case 2: Disrupted nodes are assigned IID local data
In this configuration, disrupted nodes are assigned data in an IID fashion with respect to surviving nodes. In practice, this means that we assign exactly the same number of data samples per class to all nodes, regardless of whether they survive or don’t survive the disruption.
Case 3: Disrupted nodes are assigned non-IID local data
In Case 3, the available class labels are split into two groups, $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$. Data in $\mathcal{L}_{1}$ are permanently assigned in an IID fashion, meaning that all nodes are assigned the same number of samples per $\mathcal{L}_{1}$-class. Vice versa, $\mathcal{L}_{2}$-data are assigned in an unbalanced way: a group of nodes (G1) receives a low number of samples for $\mathcal{L}_{2}$-data, while the remaining nodes (G2) receive a high amount of $\mathcal{L}_{2}$-data. As a result, all nodes see an IID distribution for $\mathcal{L}_{1}$ data, while G1 nodes see much less $\mathcal{L}_{2}$-data than G2 nodes. Case 3 is further divided into two sub-scenarios: Case 3.1 and Case 3.2. In Case 3.1, group G2 is composed of the disrupted nodes: this means that a disproportionately high number of $\mathcal{L}_{2}$-data is assigned to nodes that will disappear after disruption. With this configuration, we are testing what happens when disrupted nodes contribute much more knowledge to the training process than the other nodes. This is because larger local datasets can lead to better-trained models, thus enabling more knowledge extraction. In the second scenario (Case 3.2), we instead give disproportionately more data in the $\mathcal{L}_{2}$-label group to the 10 nodes with the lowest structural hole score (we denote them with $\mathcal{V}_{p}\subset\mathcal{V}_{s}$, i.e., the group of non-disrupted nodes at the “periphery” of the network). Group G2, in this case, is composed of the nodes in $\mathcal{V}_{p}$. This configuration is intended to simulate a situation where the majority of the $\mathcal{L}_{2}$ data is located at the “edge” of the network, reflecting scenarios where data-rich nodes are not central to the overall network structure. The key difference between the two sub-scenarios lies in what happens after the disruption: in Case 3.2, some of the G2 nodes remain connected to the network after the disruption, unlike in Case 3.1 where G2 nodes are fully isolated. This distinction is illustrated in Figure 3.
Figure 3: The figure shows the BA network after disruption. The G2 nodes of Case 3.2 are highlighted in yellow. The rest of the coloring is size-related as before.
4.4 Other settings
Dataset and data distribution
@@ 4.4 Other settings
For our experiments, we employ the widely used MNIST image dataset [27]. This dataset contains a set of handwritten digits; thus, data are divided into 10 classes (for digits from 0 to 9). The collaborative task undertaken by the nodes is then a standard image classification problem. This problem is relevant to a disaster scenario, where individuals within the affected community can leverage their smartphones or other devices with built-in cameras to capture images of the affected areas, including collapsed buildings, road blockages, or other hazardous conditions. The MNIST dataset contains 60,000 training images (approximately 6,000 per class) and 10,000 test images (1,000 per class). We evaluate the impact of the disconnection of central nodes by measuring the accuracy obtained by the local models on the standard MNIST test dataset over time. Note that all classes are equally represented in the test set and that the test set is common to all nodes.
@@
Classification on the MNIST dataset is a relatively easy learning task. Thus, in order to stress test our decentralised learning process and to capture a realistic scenario [^3] We anticipate that in scenarios where nodes represent user devices collaborating on decentralised learning tasks, the local datasets will be relatively small, as users actively collect data samples themselves., we assign very few images per class to each node. In Case 1, we distribute 7 images per class to non-disrupted nodes, leading to a local training dataset of 70 images and 6,300 images overall in the network. Recall, in fact, that in Case 1 the 10 disrupted nodes do not own any local data, so the remaining 90 nodes are the only ones assigned data. In Case 2, we keep the same IID data distribution, but this time, we involve disrupted nodes. In order to get approximately the same dataset size distributed across the whole network, this time, we assign 6 images per class to each node, leading to a local training dataset of 60 images and, globally, 6,000 images in the network (given that images cannot be fractional, this is the best trade-off to make Case 1 and Case 2 comparable). In Case 3, recall that class labels are split into two sets $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$. Images belonging to $\mathcal{L}_{1}$ are distributed in an IID fashion across all nodes by simply splitting the approximately 6,000 images among all nodes (hence, each node gets around 60 images per class). In contrast, the images from $\mathcal{L}_{2}$ are distributed unevenly to emphasize specific differences between sub-scenarios. A small number of $\mathcal{L}_{2}$ images (configurations of 10, 20, or 30 images per class) are assigned to a selected group of nodes in G1, while the remaining images are allocated to G2 nodes. In both sub-scenarios, the G2 nodes are ten, thus resulting in each of these nodes having around 500 images per $\mathcal{L}_{2}$ class.
The three different scenarios and the corresponding data distributions are summarised in Table 1.
@@
Case 3.1 IID in $\mathcal{L}_{1}$, high $\mathcal{L}_{2}$ IID (60 images/class) in $\mathcal{L}_{1}$, low $\mathcal{L}_{2}$
Case 3.2 IID (60 images/class) in $\mathcal{L}_{1}$, low $\mathcal{L}_{2}$
Low SHS: IID (60 images/class) in $\mathcal{L}_{1}$, high $\mathcal{L}_{2}$
Others: IID (60 images/class) in $\mathcal{L}_{1}$, low $\mathcal{L}_{2}$
Table 1: Summary table for the data distributions in the considered scenarios. Low $\mathcal{L}_{2}$-data is tested for $\{10,20,30\}$ images/$\mathcal{L}_{2}$-class. The remaining images per $\mathcal{L}_{2}$-class are split among the nodes receiving a high number of $\mathcal{L}_{2}$-data. Acronym SHS stands for Structural Hole Score.
Decentralised learning strategy and local model architecture
@@ 4.5 Evaluation metrics
Here, we introduce the metrics used to evaluate the robustness of decentralised learning to node disruption. A direct measure of classification performance is the accuracy achieved by nodes’ local models [^4] Note that local models here are the result of progressively averaging models obtained from neighbours. Thus they embed global knowledge. on the test set, which is common to all nodes. The individual accuracy $A_{i}(t)$ for a node $i$ at communication round $t$ is defined as the ratio between the number of correct predictions with the local model at $t$ on the test set and the total number of samples in the test set. Given that disrupted nodes do not participate in the learning any more after disruption, we do not consider the accuracy they achieve in our analyses. To obtain a compact view of the performance, we also consider the average accuracy across non-disrupted nodes, which is computed as follows.
@@
The mean accuracy of the system is defined as the mean over all the surviving nodes:
$\overline{A}(t)=\sum_{i\,\in\,\mathcal{V}_{s}}\frac{1}{|\mathcal{V}_{s}|}A_{i}(t),$ (2)
where $\mathcal{V}_{s}$ denotes the set of surviving nodes and $A_{i}(t)$ the accuracy at communication round $t$ for node $i\in\mathcal{V}_{s}$.
$\overline{A}(t)$ will be used to measure the overall system performance.
When the top 10% central nodes are disrupted, the connectivity in the graph drastically decreases. As illustrated in Figure 2(b), many nodes become completely disconnected, while others form smaller connected components [^5] Recall that a connected component is defined as a subgraph where a path exists between any node pair.. Reduced connectivity results in fewer opportunities for communication and collaboration, impairing the decentralised learning process. It is anticipated that nodes losing more connections, and particularly those becoming isolated, will be affected most severely. Thus, we also define the mean accuracy for the similarly sized connected components as follows.
Definition 4.2 (Mean accuracy within same-size connected components).
The mean accuracy over connected components of size $c$ is defined as the mean over all $\mathcal{V}_{s}$ nodes belonging to connected components of the chosen size $c$.
@@
where $\mathcal{C}_{c}$ denotes the set of nodes belonging to a connected component of size $c$.
The mean accuracy over connected components will be used to measure a more fine-grained impact of the disruption on the system, accounting for localised effects. Note that we aggregate over all nodes belonging to connected components with size $c$, regardless of whether they are in the same connected component. In the case of connected components of size 1, for example, this allows us to consider all isolated nodes together.
@@ 5 Results
In this section, we discuss the impact of central nodes being cut off from the network due to a disruption, and we focus on the three scenarios, Case 1, Case 2 and Case 3, illustrated in the previous section. Recall that in Case 1 and Case 2, nodes have an IID distribution of local data, with the difference that in Case 1, highly central nodes, i.e the ones with highest structural hole score, do not have local data, and thus they contribute only by connecting elements in the graph topology. The data that these nodes hold locally in Case 2 are redistributed in Case 1 among the other nodes, preserving the IID property among them to maintain the total number of data points equal at the overall network level. Case 3 is different from the other two, as we use it to test the impact of disruption when the data distribution across nodes is non-IID. First, in Sections 5.1-5.4, we discuss and compare Case 1 and Case 2, as in both the disrupted nodes do not enjoy more knowledge than the other nodes. Then, Case 3 will be discussed separately in Section 5.5. Due to the significantly different data distributions between Cases 1-2 and Case 3, a direct quantitative comparison would not be appropriate.
@@ 5.1 Baseline performance when no disruption occurs
Figure 4 showcases what we already anticipated in Section 4.2.2: when two learning processes have differing speeds, a disruption occurring at a fixed time $t$ will encounter two systems at distinct stages of the learning process. While this scenario realistically captures the notion that faster systems are more likely to have acquired sufficient knowledge when disrupted, comparisons between the two systems for what-if analyses may be less informative. The discrepancy between the two scenarios is depicted in Table 2, which illustrates the communication round at which the mean accuracy of the system reaches the specified accuracy threshold in both cases. As expected, Case 2 reaches the accuracy threshold faster than Case 1. Therefore, as discussed in Section 4.2.2, in this study, we chose to assess the robustness of the two systems when disruptions occur at the same average accuracy level for both. The objective is to compare the consequences of disruption in systems that had achieved similar skill levels in terms of learning.
@@ 5.2 Impact of the time of disruption
Figure 5: Accuracy curves for each surviving node in Case 1 (a) and Case 2 (b). From left to right: increasing accuracy threshold condition for the disruption. The curves are coloured based on the size of the connected component to which the corresponding node belongs after the disruption. Accuracy threshold 75% is omitted for ease of visualisation, as it only showed an intermediate behaviour between threshold 70% and 80%.
@@
We now start investigating the impact of disruptions. In Figure 5, we show the accuracy of all the surviving (i.e., non-switched-off) nodes in Case 1 and Case 2 when the disruption happens at different accuracy thresholds. The colouring here is defined based on the size of the connected component the nodes belong to after the disruption occurs, as discussed in Section 4.2.1.
@@
First, we focus on the effect of different disruption times in Figure 5. The later the disruption happens, i.e., going from left to right, the higher the mean accuracy of the system, as more and more nodes have been able to reach a higher level of accuracy before the cut-off, resulting in a narrower curve beam. This result holds for both scenarios and follows naturally after noticing that the central nodes are able to connect the network and, in Case 2, share their information for more communication rounds before getting switched off as the accuracy threshold increases. In all cases, after the disruption occurs, most curves stop improving and tend to cluster into different groups of similarly performing nodes. As the curve colours show, these groups reflect the connected component size to which the nodes belong. Recall that since the connected components’ sizes are calculated after the removal of the special nodes, nodes having connected component size 1 are isolated nodes. As we can see in Figure 5, isolated nodes have the lowest values of accuracies, and the accuracy generally increases with the increasing value of the connected component size. This comes naturally from the fact that more connected nodes enjoy more information flow due to collaborative learning, thus increasing their performance. Furthermore, by examining the curves for the two largest connected components (size 25 and size 29), we can observe that the larger the connected component size, the more a node’s performance becomes independent of the disruption time. This indicates that the connected component has sufficient data to compensate for the disruption’s effects, no matter when it occurs. Conversely, we notice a strong correlation between the accuracy level and the timing of disruptions for the isolated nodes: the later the disruption occurs, the better the accuracy level that the isolated nodes can achieve. Naturally, this is because they have longer access to more information. Connected components of intermediate sizes fall within these two boundary conditions.
@@
Interestingly, in Figure 5, we note that for later disruption times, isolated nodes, as well as nodes belonging to small connected components, reduce the accuracy they achieve with respect to the one they had achieved at the disruption time. This effect is not present for nodes in larger connected components. The intuition is that large connected components can compensate for the lack of connectivity thanks to the overall set of data residing at the nodes in the connected component.
Take-home: If the connected component is large enough, the effect of disruption can be compensated reasonably well. Otherwise, nodes in smaller connected components are not able to recover knowledge that is lost as a side effect of disruption. We analyse this effect in more detail in the next section.
@@ 5.3 Connectivity loss vs connectivity and data loss
Figure 7: Mean local accuracy difference for the largest connected component (left) and the isolated nodes (right).
@@
This is further confirmed at a more granular level when we compute the same difference but focusing on the largest connected component (Figure 7(a)) and the isolated nodes (Figure 7(b)), thus computing the difference between the within-component accuracy, as per Equation 3. The behaviour of the large connected component is similar to that observed for the system overall. Specifically, we note that (i) the condition of Case 1 is more favourable than that of Case 2 and that (2) after sufficient time, the accuracy threshold at which disruption occurred does not differentiate the two conditions. Moreover, the case of isolated nodes shows that when disruption occurs at a (relatively) low accuracy, the presence of even a small amount of additional data locally allows the model of Case 1 to reach a higher accuracy than the model trained in Case 2. However, this difference vanishes when disruption occurs at a high accuracy threshold.
Take-home: All in all, our results suggest that having more data available “somewhere” in the network (as in Case 1) is more beneficial than concentrating the same amount of data on central nodes, if those nodes are at risk of disruption (as in Case 2), while the main advantage of the latter configuration is limited to a speed-up of the decentralised learning process if no disruption occurs (as shown by Figure 4). This confirms that, as long as data are present in the network, decentralised learning is able to exploit them through collaborative training, even in the presence of significant disruptions to the network structure.